kristiannnnnt:

A number is chosen at random from the first 100 positive integers. Find the probability that the number is divisible by 5. 1/10 1/5 1/4

XxSuperxX:

what does 5/100 reduce to

XxSuperxX:

it reduces to 1/20

KingPizza:

well first, do you know all the numbers that are divisible by 5? Like 5 and 10, do you know the other numbers that are divisible by 5?

XxSuperxX:

it reduces to 1/20 so you would have 20% which = 1/5

KingPizza:

actually that's not the way to solve it, we have to use all the numbers divisible by 5 so we have to figure out how many numbers between 1 and 100 are divisible by 5

KingPizza:

@xxsuperxx wrote:
it reduces to 1/20 so you would have 20% which = 1/5
oh ok you are correct

KingPizza:

so yeah, find out how many numbers are divisible by 5, then divide with that value as the numerator and 100 as the denominator, then write it as a percentage and decimal and you'll have your answer

KingPizza:

and convert it to a fraction

Vocaloid:

To find out how many multiples of 5 there are between 1-100, divide 100/5 From there, divide by 100 again to get the probability out of 100. Then reduce.

XxSuperxX:

like you do 5/100 then you get 1/20 which then you would do 20/100 and that would equal 1/5 which also ='s 20%

XxSuperxX:

so 1/5 would be your answer